OpenStudy (anonymous):

-3x6y4 ---------- x-3y-2 Which of the following is the simplest form of this expression? Select one of the options below as your answer: A. -3x^-2y-^2B. -3x^2y^2 C. -3x^9y^-2 D. -3x^9y^6

OpenStudy (anonymous):

D

OpenStudy (anonymous):

-3 x^6 / x ^ -3 = -3x^(6--3) = -3 x^9 y^4 / y^-2 = y^(4--2) = y^6